OpenStudy (anonymous):

f(x) = 2^x for x = 3. 6 8 4 9

OpenStudy (anonymous):

For every X use 3. What is 2 cubed? 2*2*2= ?

OpenStudy (anonymous):

the calculator im using says 6

OpenStudy (anonymous):

Not 2+2+2 But rather 2*2*2=4*2=?

OpenStudy (anonymous):

\[f(3)=2^{(3)}=2*2*2=4*2=?\]

OpenStudy (anonymous):

now im really confused I was coming up with 6 on the calculator and its 8?

OpenStudy (anonymous):

Yes 2*2*2=(2*2)*2=4*2=8

OpenStudy (anonymous):

You mostprobably added instead of multiplying 2+2+2=6 2*2*2=8

OpenStudy (anonymous):

Ok now that makes sense that I typed in the calculator what you are saying and it says 8 thankyou so much for your help :)